# Sourav Banerjee and Neel Kanth Kundu

### v1

@@ I Introduction
Conventional radar systems use a dipole antenna to detect an incoming electromagnetic (EM) echo signal that induces a current signal in the receiver circuit. Their receivers are based on the bulk properties of quantum mechanics, where the EM echo signal induces the flow of a large number of electrons in the receiver antenna circuit. Recently, Rydberg atom-based RF sensors have emerged as a promising technology that is fundamentally different from the conventional dipole antenna-based RF detectors [anderson2020rydberg, gordon2019weak].
Most of the current works on Rydberg atom-based RF sensing have demonstrated the feasibility of measuring the amplitude, phase, polarization, and power of the electric field in a laboratory setup. Rydberg atom-based communications systems capable of detecting QPSK, AM, and FM modulation have also been experimentally demonstrated [meyer2018digital, borowka2022sensitivity]. However, the system design and working principle of a Rydberg atom-based radar system have not yet been investigated in the literature. Prior works on quantum radars require microwave-entangled quantum states, which require dilution refrigerators [sorelli2021detecting]. Furthermore, the joint detection at the receiver using the stored idler and reflected signal is practically challenging to implement [sorelli2021detecting]. In order to overcome these challenges, we propose a quantum radar system that uses a conventional microwave source at the transmitter. However, the detector at the receiver is replaced with the Rydberg-atom-based RF sensors.
@@
    We analyze the SNR characteristics of the Rydberg atom-based radar system, and evaluate the target velocity estimation performance by employing invariant function based frequency estimation methods [candan2021frequency].
@@ II Measurement of an Electric Field by Rydberg Atom-Based RF Sensor
Rydberg atom-based electric field sensors are composed of alkali atoms (e.g., Cesium and Rubidium) in a glass cell of dimension a few centimeters in size [anderson2020rydberg]. We consider a ladder-type four-level atom with states $|1\rangle$ (ground state), $|2\rangle$ (excited state), $|3\rangle$ and $|4\rangle$ (Rydberg states). A weak probe laser (frequency $\omega_{p}$) drives the $|1\rangle\leftrightarrow|2\rangle$ transition with Rabi frequency $\Omega_{p}$, while a strong coupling laser (frequency $\omega_{c}$) drives $|2\rangle\leftrightarrow|3\rangle$ with Rabi frequency $\Omega_{c}$. An RF field of frequency $f_{\rm RF}$ couples the Rydberg pair $|3\rangle\leftrightarrow|4\rangle$ with RF Rabi frequency $\Omega_{\rm RF}$. The Rabi frequency for a field $E$ coupling levels $i$ and $j$ is $\Omega_{ij}=\wp_{ij}/E\hbar$, where $\wp_{ij}$ is the transition dipole matrix element. The transmitted probe power through a vapor cell (adiabatic approximation) is
@@
where $N_{0}$ is the atomic density [FleischhauerEIT]. From [chen2025harnessingrydbergatomicreceivers] we get, under resonant conditions ($\Delta_{p}=\Delta_{RF}=0$, $\gamma_{41}=\gamma_{31}=0$), the imaginary part of $\rho_{21}$ is:
@@
For the sake of completeness, we need to provide a detailed discussion of the measurement of an incident RF signal using a local oscillator (LO). But for brevity, we refer to section III B of [chen2025harnessingrydbergatomicreceivers], a tabulated summary is provided in Table I.
TABLE I: Summary of LO-dressed Rydberg atomic receiver model (Section III.B of [chen2025harnessingrydbergatomicreceivers]).
@@ III System Model
In this work, we use the atomic superheterodyne detection principle proposed in [jing2020atomic, gordon2019weak]. For the radar application, the LO signal is derived from the RF source that is used for illuminating the target as shown in Fig. 2. The signal produced by the RF source which is used to illuminate the target as well as the LO is given by
@@
The variation in $E_{s}(t)$ leads to a variation in the power of the probe laser [wu2023linear]
@@
where $P_{0}$ is the DC value of the optical power and $C$ (in $\text{W}\cdot\text{m}/\text{V}$) is a constant depending on the laser frequencies and the atomic properties of the Rydberg atom [jing2020atomic, wu2023linear]. To derive the expression for the quantity $C$, we equate (9) with RF Rabi frequency equation in Table I and utilize a series of interdependent equations and parameter definitions provided in the work [chen2025harnessingrydbergatomicreceivers]. Finally, we summarize those expressions, which collectively define the dependencies leading to the final form of $C$ in Table II.
TABLE II: List of equations from [chen2025harnessingrydbergatomicreceivers], required to derive $C$ in (9).
@@
where $\mathcal{R}$ is the responsivity (in A/W), $M$ is the avalanche multiplication gain and $z(t)\sim\mathcal{N}(0,\sigma_{z}^{2})$ is the additive Gaussian noise. The noise $z(t)$ is composed of two independent noise sources: quantum shot noise arising from the discrete nature of the photons and the electronic thermal noise. The noise variance is given by [keiser2000optical]
@@ IV-B Target Velocity Estimation
Thus, the velocity of the target can be estimated by estimating the frequency $\omega$ of the output signal from the avalanche photodiode. Frequency estimation of a real sinusoidal signal has been well investigated in the literature. For example, the invariant function approach proposed in [candan2021frequency] can be used to estimate $\omega$.
To characterize the fundamental performance limit of any unbiased estimator, the well-known Cramér–Rao lower Bound (CRB) is used. The Asymptotic (ACRB) for frequency estimation is given by [candan2021frequency]
@@ V Numerical Results
Now, we compare the SNR of the standard radar with that of the Rydberg atom-based radar. The nominal values of the simulation parameters are: $P_{t}=10$ dBW, $G_{t}=10$ dB, $\sigma=1\;\text{m}^{2}$, $A_{e}=A_{s}=1\;\text{cm}^{2}$, $T_{s}=1000$K, $B_{e}=1$ MHz, $\mathcal{R}=0.6$ A/W, $M=50$, $R_{l}=1\;\text{K}\Omega$, $P_{0}=10\;\mu\text{W}$, $T=300$ K and $I_{d}=1\;\text{nA}$. By evaluating the expressions presented in Table II, and using the same parameter values as those employed in [chen2025harnessingrydbergatomicreceivers] for their computation, the value of $C$ is determined to be $6.59\times 10^{-4}\,\mathrm{W/(V/m)}$. The SNR comparison of the Rydberg atom-based radar and the standard radar is shown in Fig. 3(a). It can be observed that the SNR of the Rydberg atom radar is significantly higher ($\sim$ 40dB) than that of the conventional radar.
Next, we compare the performance of velocity estimation of the target for the Rydberg radar with that of conventional radar. We assume a target velocity of $v=100$ m/s and a sampling frequency of $f_{s}=60$ KHz, the number of samples is $N=2048$. Fig. 3(b) shows the root mean square error (RMSE) of velocity estimation obtained using the invariant function-based frequency estimation [candan2021frequency]for the Rydberg and classical radar. The plots also show the ACRBs for the velocity estimation for benchmarking. It can be observed that the RMSE of the Rydberg radar is lower than that of the classical radar. Furthermore, we observe that when employing the Doppler frequency estimation method described in [candan2021frequency], the RMSE of the Rydberg quantum radar closely follows the associated ACRB up to a larger target range, approximately 2000 meters. In contrast, the RMSE for the classical radar approaches the ACRB only at shorter distances. These observations highlight the superior performance of the Rydberg quantum radar over classical radar in terms of target velocity estimation accuracy.

### v2

@@ I Introduction
Conventional radar systems use a dipole antenna to detect an incoming electromagnetic (EM) echo signal that induces a current signal in the receiver circuit. Their receivers are based on the bulk properties of quantum mechanics, where the EM echo signal induces the flow of a large number of electrons in the receiver antenna circuit. Recently, Rydberg atom-based RF sensors have emerged as a promising technology that is fundamentally different from the conventional dipole antenna-based RF detectors [1, 6].
Most of the current works on Rydberg atom-based RF sensing have demonstrated the feasibility of measuring the amplitude, phase, polarization, and power of the electric field in a laboratory setup. Rydberg atom-based communications systems capable of detecting QPSK, AM, and FM modulation have also been experimentally demonstrated [9, 2]. However, the system design and working principle of a Rydberg atom-based radar system have not yet been investigated in the literature. Prior works on quantum radars require microwave-entangled quantum states, which require dilution refrigerators [11]. Furthermore, the joint detection at the receiver using the stored idler and reflected signal is practically challenging to implement [11]. In order to overcome these challenges, we propose a quantum radar system that uses a conventional microwave source at the transmitter. However, the detector at the receiver is replaced with a Rydberg-atom-based RF sensor.
@@
    We analyze the SNR characteristics of the Rydberg atom-based radar system, and evaluate the target velocity estimation performance by employing invariant function based frequency estimation methods [3].
@@ II Measurement of an Electric Field by Rydberg Atom-Based RF Sensor
Rydberg atom-based electric field sensors are composed of alkali atoms (e.g., Cesium and Rubidium) in a glass cell of dimension a few centimeters in size [1]. We consider a ladder-type four-level atom with states $|1\rangle$ (ground state), $|2\rangle$ (excited state), $|3\rangle$ and $|4\rangle$ (Rydberg states). A weak probe laser (frequency $\omega_{p}$) drives the $|1\rangle\leftrightarrow|2\rangle$ transition with Rabi frequency $\Omega_{p}$, while a strong coupling laser (frequency $\omega_{c}$) drives $|2\rangle\leftrightarrow|3\rangle$ with Rabi frequency $\Omega_{c}$. An RF field of frequency $f_{\rm RF}$ couples the Rydberg pair $|3\rangle\leftrightarrow|4\rangle$ with RF Rabi frequency $\Omega_{\rm RF}$. The Rabi frequency for a field $E$ coupling levels $i$ and $j$ is $\Omega_{ij}=\wp_{ij}E/\hbar$, where $\wp_{ij}$ is the transition dipole matrix element. The transmitted probe power through a vapor cell (adiabatic approximation) is
@@
where $N_{0}$ is the atomic density [5]. From [4] we get, under resonant conditions ($\Delta_{p}=\Delta_{RF}=0$, $\gamma_{41}=\gamma_{31}=0$), the imaginary part of $\rho_{21}$ is:
@@
For the sake of completeness, we need to provide a detailed discussion of the measurement of an incident RF signal using a local oscillator (LO). But for brevity, we refer to section III B of [4], a tabulated summary is provided in Table I.
TABLE I: Summary of LO-dressed Rydberg atomic receiver model (Section III.B of [4]).
@@ III System Model
In this work, we use the atomic superheterodyne detection principle proposed in [7, 6]. For the radar application, the LO signal is derived from the RF source that is used for illuminating the target as shown in Fig. 2. The signal produced by the RF source which is used to illuminate the target as well as the LO is given by
@@
The variation in $E_{s}(t)$ leads to a variation in the power of the probe laser [12]
@@
where $P_{0}$ is the DC value of the optical power and $C$ (in $\text{W}\cdot\text{m}/\text{V}$) is a constant depending on the laser frequencies and the atomic properties of the Rydberg atom [7, 12]. To derive the expression for the quantity $C$, we equate (9) with RF Rabi frequency equation in Table I and utilize a series of interdependent equations and parameter definitions provided in the work [4]. Finally, we summarize those expressions, which collectively define the dependencies leading to the final form of $C$ in Table II.
TABLE II: List of equations from [4], required to derive $C$ in (9).
@@
where $\mathcal{R}$ is the responsivity (in A/W), $M$ is the avalanche multiplication gain and $z(t)\sim\mathcal{N}(0,\sigma_{z}^{2})$ is the additive Gaussian noise. The noise $z(t)$ is composed of two independent noise sources: quantum shot noise arising from the discrete nature of the photons and the electronic thermal noise. The noise variance is given by [8]
@@ IV-B Target Velocity Estimation
Thus, the velocity of the target can be estimated by estimating the frequency $\omega$ of the output signal from the avalanche photodiode. Frequency estimation of a real sinusoidal signal has been well investigated in the literature. For example, the invariant function approach proposed in [3] can be used to estimate $\omega$.
To characterize the fundamental performance limit of any unbiased estimator, the well-known Cramér–Rao lower Bound (CRB) is used. The Asymptotic (ACRB) for frequency estimation is given by [3]
@@ V Numerical Results
Now, we compare the SNR of the standard radar with that of the Rydberg atom-based radar. The nominal values of the simulation parameters are: $P_{t}=10$ dBW, $G_{t}=10$ dB, $\sigma=1\;\text{m}^{2}$, $A_{e}=A_{s}=1\;\text{cm}^{2}$, $T_{s}=1000$K, $B_{e}=1$ MHz, $\mathcal{R}=0.6$ A/W, $M=50$, $R_{l}=1\;\text{K}\Omega$, $P_{0}=10\;\mu\text{W}$, $T=300$ K and $I_{d}=1\;\text{nA}$. By evaluating the expressions presented in Table II, and using the same parameter values as those employed in [4] for their computation, the value of $C$ is determined to be $6.59\times 10^{-4}\,\mathrm{W/(V/m)}$. The SNR comparison of the Rydberg atom-based radar and the standard radar is shown in Fig. 3(a). It can be observed that the SNR of the Rydberg atom radar is significantly higher ($\sim$ 40dB) than that of the conventional radar.
Next, we compare the performance of velocity estimation of the target for the Rydberg radar with that of conventional radar. We assume a target velocity of $v=100$ m/s and a sampling frequency of $f_{s}=60$ KHz, the number of samples is $N=2048$. Fig. 3(b) shows the root mean square error (RMSE) of velocity estimation obtained using the invariant function-based frequency estimation [3]for the Rydberg and classical radar. The plots also show the ACRBs for the velocity estimation for benchmarking. It can be observed that the RMSE of the Rydberg radar is lower than that of the classical radar. Furthermore, we observe that when employing the Doppler frequency estimation method described in [3], the RMSE of the Rydberg quantum radar closely follows the associated ACRB up to a larger target range, approximately 2000 meters. In contrast, the RMSE for the classical radar approaches the ACRB only at shorter distances. These observations highlight the superior performance of the Rydberg quantum radar over classical radar in terms of target velocity estimation accuracy.